# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

### v2

@@ Abstract
Although many popular reinforcement learning algorithms are underpinned by $f$-divergence regularization, their sample complexity with respect to the regularized objective still lacks a tight characterization. In this paper, we analyze $f$-divergence-regularized offline policy learning. For reverse Kullback–Leibler (KL) divergence, arguably the most commonly used one, we give the first $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for contextual bandits, surpassing existing $\widetilde{O}(\epsilon^{-1})$ bound under all-policy concentrability and $\widetilde{O}(\epsilon^{-2})$ bound under single-policy concentrability. Our analysis for general function approximation leverages the principle of pessimism in the face of uncertainty to refine a mean-value-type argument to its extreme. This in turn leads to a novel moment-based technique, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class. We further propose a lower bound, demonstrating that a multiplicative dependency on single-policy concentrability is necessary to maximally exploit the strong convexity of reverse KL. In addition, for $f$-divergences with strongly convex $f$, to which reverse KL does not belong, we show that the sharp sample complexity $\widetilde{\Theta}(\epsilon^{-1})$ is achievable even without single-policy concentrability. In this case, the algorithm design can get rid of pessimistic estimators. We further extend our analysis to dueling bandits, and we believe these results take a significant step toward a comprehensive understanding of $f$-divergence-regularized policy learning.
@@ 1 Introduction
where $r$ is the reward, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature. When $\pi^{\mathsf{ref}}$ is uniform, 1.1 reduces to the entropy-regularized objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Levine and Koltun, 2013; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024). KL regularization has also been widely used in the RL fine-tuning of large language models (Ouyang et al., 2022; Rafailov et al., 2023), where $\pi^{\mathsf{ref}}$ is the base model. Given its widespread use, there has been a surge of interest in understanding the role of KL regularization in RL by both empirical studies (Ahmed et al., 2019; Liu et al., 2019) and theoretical analysis (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022). There are also lines of research on KL regularization in online learning (Cai et al., 2020; He et al., 2022; Ji et al., 2023) and convex optimization (Neu et al., 2017). However, the performance metric in most of these studies is still the unregularized reward maximization objective, under which the sample complexity is at least $\Omega(\epsilon^{-2})$. (See Section A.1 for detailed discussions)
Table 1: Comparison of sample complexity bounds for finding $\epsilon$-optimal policy for offline contextual bandits with KL- and (strongly convex) $f$-divergence regularization. Constants and $\mathrm{polylog}$ factors are omitted here except the metric entropy $\log\mathcal{N}$. “Reverse-KL” stands for KL-regularized contextual bandits and “$f$-divergence w/ s.c., $f$” for the counterpart with an $\alpha$-strongly convex $f$. The two existing upper bounds are adapted from the implicit form in Xiong et al. (2024, Theorem 3.1) and Zhao et al. (2024, Theorem 3.3 and Theorem 4.4.), of which the detailed adaptions are deferred to Appendix A. Also see Remark 2.8 for discussion on the relationship between $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$
@@
Several recent papers (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024; Foster et al., 2025) switched the focus to analyzing the suboptimality defined via the regularized objective (1.1), under which an $\Omega(\epsilon^{-1})$ sample complexity is possible (Zhao et al., 2024, Theorem 3.6). However, even restricted to the pure $\mathrm{i.i.d.}$ setting, existing analyses in this vein either result in still $\widetilde{O}(\epsilon^{-2})$ bounds (Xiong et al., 2024; Xie et al., 2024) or has a stringent all-policy concentrabiity dependency in their upper bounds (Zhao et al., 2024) (See Section 2.1 for details on coverage conditions). Thus, there are by far no perfectly matching bounds down to both the dependency of $\epsilon^{-1}$ and concrete coverage conditions for KL-regularized offline decision making. Moreover, these theoretical analyses set KL as the right target by default; but reverse KL is the $f$-divergence with $f(x)=x\log x$, which is merely convex. Therefore, it is also unknown whether those choices with even nicer (e.g., strongly convex) $f$, whose performance against the reward maximization objective can be provably promising (Zhan et al., 2022; Gabbianelli et al., 2024; Huang et al., 2024), can enjoy a better coverage dependency in the sample complexity under the corresponding regularized objectives. These perspectives motivate an open problem:
What is the weakest coverage condition for offline policy learning to be minimax optimal with respect to $f$-divergence-regularized objectives?
We are the first to concretely investigate this question by proposing sharp sample complexity analyses for two representative ad subclasses of $f$-divergence. In particular, for contextual bandits with KL regularization, we present sample complexity guarantee with linearly dependence on single-policy coverage ratio. Our lower bound further indicates that a multiplicative dependency on single-policy concentrability is necessary. Surprisingly, for $f$-divergence with $\alpha$-strongly-convex $f$, we prove nearly matching sample complexity bounds of $\widetilde{\Theta}(\alpha^{-1}\eta\epsilon^{-1})$ for the first time, settling the dependence on coverage completely. For the ease of comparison, we adapt previous sample complexity bounds to our setting and notation, and summarize them in Table 1.
@@ 1.1 Contributions
  • For arguably the most commonly used KL regularization, we are the first to employ the principle of pessimism (Jin et al., 2021) to achieve a nearly tight sample complexity under single-policy concentrability. Our analysis roadmap leverages the strong convexity of KL and pessimism of the reward estimator, to refine a mean-value-type risk upper bound to its extreme which in turn leads to a novel moment-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
  • Under KL-regularization, we prove a lower bound growing linearly the density-ratio-based single-policy concentrability, which also strictly improved upon previous works (Zhao et al., 2024; Foster et al., 2025) in offline setting. This lower bound manifests the necessity of a multiplicative dependency on single-policy concentrability and firstly closes the gap between upper and lower bounds under a technically nontrivial case.
@@
  • For any $f$-divergence with strongly convex $f$, we are able to design a truly lightweight algorithm free of pessimism-based gadgets and still obtain the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity certified by a matching lower bound without coverage conditions.
@@
  • We extend all algorithmic ideas and hard instance constructions above to $f$-divergence-regularized contextual dueling bandits, achieving $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity bounds. Moreover, all algorithms work for general reward function classes with finite metric entropy.
1.2 Additional Related Work
We review two additional lines of theoretical progress that are closely related to our algorithm design and analysis.
Pessimism in offline RL.
@@ 1.2 Additional Related Work
The principle of pessimism has been underpinning offline RL for both the tabular (Rashidinejad et al., 2021) and function approximation (Jin et al., 2021) settings under the name of lower confidence bound (LCB). For contextual bandits, it is behind the adaptively optimal sample complexity analysis (Li et al., 2022). Shi et al. (2022) proposed a LCB-based model-free algorithm for tabular RL with near-optimal guarantee. Jin et al. (2021); Xiong et al. (2022); Di et al. (2023) utilized LCB in conjunction with the classic least-square value iteration paradigm to derive $\widetilde{O}(\epsilon^{-2})$ sample complexity results for model-free RL with function approximation. The line of work from Rashidinejad et al. (2021); Xie et al. (2021b) to Li et al. (2024) settled the sample complexity of tabular model-based RL via pessimistic estimators exploiting the variance information. It is also possible to leverage the idea of pessimism to design model-based algorithms under general function approximation that are at least statistically efficient (Xie et al., 2021a; Uehara and Sun, 2021; Wang et al., 2024).
@@
However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting. The algorithmic idea of using pessimistic least-square estimators under general function approximation in Jin et al. (2021); Di et al. (2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, which cannot lead to the desired sample complexity of our objective by direct adaption.
Offline contextual dueling bandits.
@@
CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014). Since the empirical breakthrough of preference-based RL fine-tuning of LLMs (Ouyang et al., 2022), the theory of offline CDBs has received more attention under linear (Zhu et al., 2023; Xiong et al., 2024) and general (Zhan et al., 2022; Zhao et al., 2024; Song et al., 2024; Huang et al., 2024) function approximation. Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, viz., our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
@@
The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout the paper. For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$, and $y_{n}=\Theta(x_{n})$ if $x_{n}=O(y_{n})$ and $x_{n}=\Omega(y_{n})$. We further employ $\widetilde{O}(\cdot),\widetilde{\Omega}(\cdot)$, and $\widetilde{\Theta}$ to hide $\mathrm{polylog}$ factors. For countable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$. For $g:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|g\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|g(x)|$. For a pair of probability measures $P\ll Q$ on the same space and function $f:\mathbb{R}_{+}\to\mathbb{R}$, their $f$-divergence is $D_{f}(P\|Q)\coloneqq\int f({\,\mathrm{d}}P/{\,\mathrm{d}}Q){\,\mathrm{d}}Q$. Specifically, when $f(x)=x\log x$, $f$-divergence becomes KL divergence denoted as ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and when $f(x)=|x-1|/2$, it becomes the total variation (TV) distance, which is denoted as ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$.
@@ 2 KL-regularized Contextual Bandits
In this section, we propose a pessimism-based algorithm, dubbed PCB-KL, for offline KL-regularized contextual bandits. In the following subsections, we also showcase our key technical novelty which couples the pessimism of our reward estimator to the non-trivial curvature property of KL regularization.
@@ 2.1 Problem Setup
$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}r(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$ (2.1)
where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization. For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$. This type of “behavior regularization” has been also been studied in Zhan et al. (2022). The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is given by (See, e.g., Zhang 2023, Proposition 7.16)
@@
A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$. To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains the estimator $\widehat{r}$. More specifically, we work with general function approximation under realizability, which is as follows.
@@
We also need a standard condition to control the complexity of ${\mathcal{G}}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
@@
When ${\mathcal{G}}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of ${\mathcal{G}}$ is given by $\mathcal{N}_{{\mathcal{G}}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{{\mathcal{G}}}$ is mild.
Concentrability.
The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024), which quantifies the ability of the behavioral policy to generate diverse actions. We first define the density-ratio-based concentrability as follows.
@@
In the definition above, small all-policy concentrability intuitively corresponds to $\text{supp}(\pi^{\mathsf{ref}})$ covering all possible inputs. On the other hand, small single-policy concentrability means that $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$. In this paper, in addition to density-ratio-based concentrability, we also adopt the following $D^{2}$-based concentrabilites to better capturing the nature of function class ${\mathcal{G}}$. In detail, we first introduce $D^{2}$-divergence as follows.
@@
The “eluder dimension”-type Definition 2.5 is directly inspired by Di et al. (2023); Zhao et al. (2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $g$ and $h$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$. Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair. We are now ready to define the two notions of concentrability conditions.
@@
Assumption 2.7 indicates that the errors on the distributions of state-action pairs $\rho\times\pi^{*}$ can be bounded by the error on the samples from $\rho\times\pi^{\mathsf{ref}}$ up to some constant. For both types, the single-policy concentrability assumption is strictly weaker than the all-policy concentrability assumption. However, in general, the two quantities characterizing single-policy concentrability $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to constant factors.
Remark 2.8.
In particular, we have $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$, indicating that $C^{\pi^{*}}$ subsumes $D^{2}_{\pi^{*}}$ when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ can be seen as constant.

### v3

@@ Abstract
Many offline reinforcement learning algorithms are underpinned by $f$-divergence regularization, but their sample complexity defined with respect to regularized objectives still lacks tight analyses, especially in terms of concrete data coverage conditions. In this paper, we study the exact concentrability requirements to achieve the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity for offline $f$-divergence-regularized contextual bandits. For reverse Kullback–Leibler (KL) divergence, arguably the most commonly used one, we achieve an $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for the first time via a novel pessimism-based analysis, surpassing existing $\widetilde{O}(\epsilon^{-1})$ bound under all-policy concentrability and $\widetilde{O}(\epsilon^{-2})$ bound under single-policy concentrability. We also propose a near-matching lower bound, demonstrating that a multiplicative dependency on single-policy concentrability is necessary to maximally exploit the curvature property of reverse KL. Moreover, for $f$-divergences with strongly convex $f$, to which reverse KL does not belong, we show that the sharp sample complexity $\widetilde{\Theta}(\epsilon^{-1})$ is achievable even without pessimistic estimation or single-policy concentrability. We further corroborate our theoretical insights with numerical experiments and extend our analysis to contextual dueling bandits. We believe these results take a significant step towards a comprehensive understanding of objectives with $f$-divergence regularization.
@@ 1 Introduction
where $r$ is the reward, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature. When $\pi^{\mathsf{ref}}$ is uniform, 1.1 reduces to the entropy-regularized objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Levine and Koltun, 2013; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024). KL regularization has also been widely used in the RL fine-tuning of large language models (Ouyang et al., 2022; Rafailov et al., 2023), where $\pi^{\mathsf{ref}}$ is the base model. Given its widespread use, there has been a surge of interest in understanding the role of KL regularization in RL by both empirical studies (Ahmed et al., 2019; Liu et al., 2019) and theoretical analysis (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022). There are also lines of research on KL regularization in online learning (Cai et al., 2020; He et al., 2022; Ji et al., 2023) and convex optimization (Neu et al., 2017). However, most of these works still study the unregularized reward maximization objective, against which the sample complexity is at least $\Omega(\epsilon^{-2})$. [^1] See Section A.1 for detailed reasons.
Table 1: Comparison of sample complexity bounds for finding $\epsilon$-optimal policy for offline contextual bandits with KL- and (strongly convex) $f$-divergence regularization. Constants and $\mathrm{polylog}$ factors are omitted here except the metric entropy $\log\mathcal{N}$. “Reverse-KL” stands for KL-regularized contextual bandits and “$f$-divergence w/ s.c., ff” for the counterpart with an $\alpha$-strongly convex $f$. The two existing upper bounds are adapted from the implicit form in Xiong et al. (2024, Theorem 3.1) and Zhao et al. (2024, Theorem 3.3 and Theorem 4.4), of which the detailed adaptions are deferred to Appendix A. The relationship between $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$ is detailed in Section 2.1.
@@
Several recent papers (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024; Foster et al., 2025; Aminian et al., 2025) switched the focus to analyzing the sub-optimality defined via the regularized objective (1.1), under which an $\Omega(\epsilon^{-1})$ sample complexity is possible (Zhao et al., 2024, Theorem 3.6). However, even restricted to the pure $\mathrm{i.i.d.}$ setting, existing analyses in this vein either result in still $\widetilde{O}(\epsilon^{-2})$ bounds (Xiong et al., 2024; Xie et al., 2024) or has stringent (local) all-policy concentrability dependencies in their upper bounds (Zhao et al., 2024; Aminian et al., 2025). [^2] See Section 2.1 for details on coverage conditions. Thus, there are by far no tight bounds in terms of both the dependency of $\epsilon^{-1}$ and data coverage conditions for KL-regularized offline decision making. In addition, all analyses above set KL as the right target by default; but reverse KL is the $f$-divergence with $f(x)=x\log x$, which is merely convex. Therefore, it is also unknown whether $f$-divergence regularizers with even nicer (e.g., strongly convex) $f$, whose performance against the reward maximization objective are provably promising (Zhan et al., 2022; Gabbianelli et al., 2024; Huang et al., 2025b), can enjoy a better coverage dependency in their sample complexity when the corresponding regularized objectives serve as the performance metric. Because data coverage (i.e., concentrability) conditions captures the crucial distributional shift issue in offline RL (Levine et al., 2020), the aforementioned perspectives motivate a pivotal open problem:
What is the weakest coverage condition required for offline learning to be near-optimal with respect to $f$-divergence-regularized objectives?
We attack this problem by showing near-optimal sample complexity with matching concentrability dependencies for two representative subclasses of $f$-divergence. First, for contextual bandits with KL regularization, we achieve a near-optimal sample complexity guarantee with linear dependence on single-policy coverage ratio. Our novel lower bound further indicates that this multiplicative dependency on single-policy concentrability is necessary. Surprisingly, for $f$-divergence with $\alpha$-strongly-convex $f$, we prove nearly matching sample complexity bounds of $\widetilde{\Theta}(\alpha^{-1}\eta\epsilon^{-1})$, eliminating the dependence on coverage for the first time. For the ease of comparison, we adapt existing counterparts under our notation to the offline setting and summarize them in Table 1.
@@ 1.1 Contributions
  • For KL regularization, we propose a pessimism-based algorithm achieving the tight sample complexity under single-policy concentrability. We also obtain a lower bound that linearly scale with the density-ratio-based single-policy concentrability. Both results strictly improves upon previous works (Zhao et al., 2024; Foster et al., 2025) in the offline setting, showing that single-policy concentrability is both sufficient and necessary to achieve the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity.
@@
  • Technically speaking, our analysis exploits the strong convexity of KL and pessimism of the reward estimator, to refine a mean-value-type risk upper bound (Lemma 2.14) to its, which in turn leads to a novel moment-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class. To the best of our knowledge, this machinery has not been used in the standard analysis of existing offline RL algorithms and may be of independent interest.
@@
  • For $f$-divergence-regularized objectives with strongly convex $f$, we design a truly lightweight algorithm free of pessimism-based gadgets and still obtain the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity certified by a matching lower bound without coverage conditions.
@@
  • We verify the statistical rates above in numerical experiments, and demonstrate the versatility of all algorithmic and constructive proof ideas above by extending them to $f$-divergence-regularized contextual dueling bandits (CDBs), achieving similar $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity bounds. Moreover, all algorithms are applicable for reward function classes with small metric entropy.
1.2 Key Related Work
@@
We review two key lines of theoretical progress that are relevant to our algorithm design and analysis. Pessimism in offline RL. The principle of pessimism has been underpinning offline RL for both the tabular (Rashidinejad et al., 2021) and function approximation (Jin et al., 2021) settings under the name of lower confidence bound (LCB). For contextual bandits, it is behind the adaptively optimal sample complexity analysis (Li et al., 2022). Shi et al. (2022) proposed a LCB-based model-free algorithm for tabular RL with near-optimal guarantee. Jin et al. (2021); Xiong et al. (2022); Di et al. (2024) utilized LCB in conjunction with the classic least-square value iteration paradigm to derive $\widetilde{O}(\epsilon^{-2})$ sample complexity results for model-free RL with function approximation. The line of work from Rashidinejad et al. (2021); Xie et al. (2021b) to Li et al. (2024) settled the sample complexity of tabular model-based RL via pessimistic estimators exploiting the variance information. It is also possible to leverage the idea of pessimism to design model-based algorithms under general function approximation that are at least statistically efficient (Xie et al., 2021a; Uehara and Sun, 2021; Wang et al., 2024). The principle of pessimism has also been applied in counterfactual empirical risk minimization (Swaminathan and Joachims, 2015; London and Sandler, 2019) and offline policy learning (Sakhi et al., 2023; 2024), which are orthogonal to our contributions.
@@
However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting. The algorithmic idea of using pessimistic least-square estimators under general function approximation in Jin et al. (2021); Di et al. (2024) is similar to ours, but their sub-optimality gap is bounded by the sum of bonuses, which cannot directly lead to the desired sample complexity of our objective.
@@
Offline CDBs. CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in the classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014). Since the empirical breakthrough of preference-based RL fine-tuning of LLMs (Ouyang et al., 2022), the theory of offline CDBs has received more attention under linear function approximation (Zhu et al., 2023; Xiong et al., 2024) and general function approximation (Zhan et al., 2022; Zhao et al., 2024; Song et al., 2024; Huang et al., 2025b). Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, namely, our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
@@
[…]
@@ 2 KL-regularized Contextual Bandits
In this section, we introduce a pessimism-based algorithm, PCB-KL, for offline KL-regularized contextual bandits. We then showcase our novel analysis techniques for PCB-KL, which couples the algorithmic pessimism with the curvature property of KL-regularized objectives.
@@ 2.1 Problem Setup
$J_{\eta}(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg[r(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg],$ (2.1)
where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization. For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$, which is similar to the type of “behavior regularization” studied in Zhan et al. (2022). The unique optimal policy $\pi^{*}_{\eta}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J_{\eta}(\pi)$ is given by (See, e.g., Zhang 2023, Proposition 7.16) [^3] We suppress $J_{\eta}$ into $J$ and $\pi^{*}_{\eta}$ into $\pi^{*}$ when they are clear in context in the following presentation.
@@
A policy $\pi$ is said to be $\epsilon$-optimal if $\mathrm{SubOpt}_{\mathrm{RKL}}(\pi)\coloneqq J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy using $\mathcal{D}$. Note that $\mathrm{SubOpt}_{\mathrm{RKL}}(\cdot)$ is defined through 2.1 and thus depends on $\eta$. To ensure that $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains an estimator $\widehat{r}$. More specifically, we work with general function approximation under realizability, which is as follows.
@@
We also employ the standard notion of covering number (Wainwright, 2019, Definition 5.1) as the complexity measure of the reward function class $\mathcal{G}$}.
@@
Assumption 2.3 allowing $\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ to be roughly negligible is arguably mild. For example, when $\mathcal{G}$} is the class of linear functions of dimension $d$ and radius $R$, the covering number is $\mathcal{N}_{{\mathcal{G}}}(\epsilon)=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), which satisfies Assumption 2.3.
Concentrability. The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024), which quantifies the ability of the behavioral policy to generate diverse actions. We first define the density-ratio-based concentrability as follows.
@@
In the definition above, small all-policy concentrability intuitively corresponds to $\text{supp}(\pi^{\mathsf{ref}})$ covering all possible inputs. On the other hand, small single-policy concentrability means that $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$. In this paper, in addition to density-ratio-based concentrability, we also adopt the following $D^{2}$-based concentrabilites to better capturing the nature of function class $\mathcal{G}$}. In detail, we start with the $D^{2}$-divergence as follows.
@@
The “eluder dimension”-type Definition 2.5 is directly inspired by Di et al. (2024); Zhao et al. (2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $g$ and $h$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$. Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair.
Remark 2.6.
For the tabular setting, a direct computation yields $D^{2}(s,a)=(\rho(s)\pi^{\text{ref}}(a|s))^{-1}$, which can be estimated by the visitation frequency empirically. Under linear function approximation, it is well known that $D^{2}(s,a)=\|\bm{\phi}(s,a)\|^{2}_{\bm{\Sigma}^{-1}}$ under mild conditions of the parameter space, where $\bm{\Sigma}=\mathbb{E}_{\rho\times\pi^{\text{ref}}}\bm{\phi}(s,a)\bm{\phi}(s,a)^{\top}$ is the covariance matrix, which can be estimated by empirical covariance matrices in practice, potentially with ridge regularization. For more general function classes like neural networks, the $D^{2}$ can also be efficiently approximated by heuristics as discussed in Xiong et al. (2024); Gupta et al. (2024); Xu et al. (2025).
We are now ready to define the two notions of concentrability conditions.
@@
Assumption 2.8 indicates that the errors on the distributions of state-action pairs $\rho\times\pi^{*}$ can be bounded by the error on the samples from $\rho\times\pi^{\mathsf{ref}}$ up to some constant. For both types, the single-policy concentrability assumption is strictly weaker than the all-policy concentrability assumption. However, in general, the two quantities characterizing single-policy concentrability $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to constant factors. In particular, we have $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$, indicating that $C^{\pi^{*}}$ subsumes $D^{2}_{\pi^{*}}$ when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ can be seen as constants. We refer the reader to Appendix B for a further discussion on the relation between $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$.